OpenStudy (anonymous):

Does the graph of function xn/(x3−4x+5) has asymptote horisontale for: a). n=3 b).n<3 can u help me please ?

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

It's *x^n

OpenStudy (amistre64):

you have a rule for polynomials, otherwise you can simply divide everything by the higest poower of x

OpenStudy (jdoe0001):

rule is if the numerator has a bigger degree than the denominator it has none if the numerator has a smaller degree than the denominator horizontal asymptote is at y = 0

OpenStudy (amistre64):

what do you recall of these rules?

OpenStudy (amistre64):

jdoe seems to know them ...

OpenStudy (anonymous):

thank you so much !!! :) @jdoe0001 @amistre64

OpenStudy (jdoe0001):

yw

OpenStudy (anonymous):

can you help me with that : lim \[\sqrt[3]{x^{2}}(\sqrt{x+1}-\sqrt{x^{2}-1})\] as x approch to infinity @jdoe0001